OpenStudy (anonymous):

Factor completely. 2ay2 + 5ay - 3a -a(2y - 1)(y + 3) a(2y - 3)(y + 1) a(2y - 1)(y + 3)

Directrix (directrix):

@Shannon20150 From the expression 2ay^2 + 5ay - 3a factor out the "a" which is common to all three terms.

Directrix (directrix):

2ay^2 + 5ay - 3a = a(2 y^2 + 5y - 3) = a * ( ? ) * ( ? )